My question: Our class average has dropped from 88 to 76, what is the percent of the decrease?

Answer by jim_thompson5910(35256)

Difference: 88-76 = 12

Divide the difference by the original amount: 12/88 = 3/22 = 0.13636 = 13.636%

The percent decrease is approximately 13.636%